# Nearly geodesic surfaces are filling

### v2

@@ 2.2 A natural metric on the Grassmann bundles
using the parallel transport. Given $\Pi\in(\mathcal{G}_{2})_{x}M$ and $\alpha:(-\eta,\eta)\rightarrow M$ such that $\alpha^{\prime}(0)\perp\Pi$ in $T_{x}M$, let $Z(t)$ be the parallel transport of $\Pi$ along $\alpha$. Let $\sigma:(-\eta,\eta)\rightarrow\mathcal{G}_{2}M$ be the curve
@@
2.3 Fundamental domains and filling
@@
By the first paragraph of [Rat19, 6.7], there exists a convex $D$. Furthermore, [Rat19, Theorem 6.7.1] proves that $\bar{D}$ is a convex polyhedron. Finally, [Rat19, Theorem 6.7.4(2), Theorem 6.7.5] prove that the fundamental domain $D$ is exact: for any side $F$ of $D$, there exists a unique $g\in\Gamma$ such that
@@
A closed hyperbolic $n$-manifold admits a fundamental domain $D\subset\mathbb{H}^{n}$ which is a convex polyhedron, whose sides are partitioned into pairs $\{\{F_{1},F_{1}^{\prime}\},\cdots,\{F_{m},F_{m}^{\prime}\}\}$, such that for each pair of sides $F_{i},F_{i}^{\prime}$, there exists a unique $g_{i}\in\Gamma$ which maps $F_{i}$ isometrically onto $F_{i}^{\prime}$. Moreover, the group $\Gamma$ is generated by $\{g_{1},\cdots,g_{m}\}$.
We call a surface $S\hookrightarrow M$ filling if MSM-S consists of contractible components. We always assume surfaces in hyperbolic $3$-manifolds are quasi-Fuchsian. Let $\pi:\mathbb{H}^{3}\rightarrow M$ be the covering projection. Then $\pi(\partial D)$ is a filling closed surface which is homotopically trivial. This motivates us to define a filling homotopy class of surfaces as a property that should hold for every representative [^1] We thank Chao Wang for clarifications regarding filling surfaces versus filling homotopy classes.. A filling homotopy class $[S]$ intersects every homotopically nontrivial loop of $M$ which means $\pi_{1}(M-S)\rightarrow\pi_{1}(M)$ is trivial. We call a disconnected space contractible if each of its components is contractible. We record the following useful lemma from Rubinstein-Sageev.
@@
Let $M$ be a closed hyperbolic $3$-manifold and $[S]$ a filling homotopy class of surfaces. Then a least area representative $S$ is filling.
Thus similar to the boundary of a fundamental domain, a least area surface in a filling homotopy class has an area lower bound by the linear isoperimetric inequality (see Theorem 1.11).
@@ 3 Strongly filling surfaces in hyperbolic $3$-manifolds
3.1 Some lemmas on intersections and filling
We first define the intersection number between $\gamma]\in\pi_{1}(M)$ and a homotopy class $[S]$ of immersed surfaces. By applying homotopies, we may assume an immersed closed curve $\gamma:S^{1}\rightarrow M$ and an immersed surface $f:S\rightarrow M$ intersect at finitely many points $x_{1},\cdots,x_{n}$, such that for each $i\in\{1,\cdots,n\}$, there exist $j\in\{1,\cdots,i(\gamma)\}$, $k\in\{1,\cdots,i(S)\},t_{ij}\in S^{1}$, and $y_{ik}\in S$ with
@@ 3.1 Some lemmas on intersections and filling
$|\gamma\cap S|=\sum_{i=1}^{n}i(\gamma)\cdot i(S)$
@@
If a Fuchsian surface is filling, it must self-intersect sufficiently to intersect all of its own geodesics transversally. A Fuchsian surface minimizes the intersection number with a closed geodesic in its homotopy class.
@@
Lift $M$ to a fundamental domain $D\subset\mathbb{H}^{3}$. The preimages of $\gamma$ and $S$ in $D$ are finitely many geodesic arcs and geodesic polygons. If an arc which is a subset of a geodesic line $\tilde{\gamma}$ has a transverse intersection with a totally geodesic plane $P$, then the circle $\partial P$ separates the two endpoints of $\tilde{\gamma}$. By the Morse lemma, any homotopy of $\gamma$ induces a homotopy of $\tilde{\gamma}$, fixing its limit points. Similarly, any homotopy of $S$ induces a homotopy of $P$, fixing the circle at infinity. Thus no homotopy decreases the intersection number between $\beta$ and $P$. Applying this argument to all transverse intersections between the preimages of $\gamma$ and the preimages of $S$ through $D$, we conclude that the minimal number is realized by taking the geodesic representatives of both $\gamma$ and $S$. ∎
We call the intersection between a Fuchsian surface and a closed geodesic in a hyperbolic manifold essential. The strong interplay between topology and geometry is a recurring theme in hyperbolic geometry. If $M=S^{2}\times S^{1}$ and $\gamma\subset S^{2}$ a geodesic great circle, then $\gamma\times S^{1}\subset M$ is a compressible totally geodesic torus. Thus 3.2 fails in general.
@@
If a Fuchsian surface $S$ intersects a geodesic $\gamma$ transversally, then for any $S^{\prime}$ homotopic to $S$, $\pi_{1}(M-S^{\prime})$ does not contain $\gamma]$. Moreover, if a Fuchsian surface $S$ is filling, its homotopy class and commensurability class are strongly filling.
@@
3.2 and 3.3 do not generalize to least area non-Fuchsian surfaces in hyperbolic manifolds: 3.22 proves every such surface has redundant intersections with some closed geodesics.
@@
Let $S\subset M$ be a filling surface and $M^{\prime}$ is a cover of $M$. Then $S$ lifts to a (possibly disconnected or noncompact) filling surface $S^{\prime}\subset M^{\prime}$.
@@
(3.1) $\mathcal{P}(M)\coloneqq\mathcal{G}_{2}M/\sim$
@@
Definition 3.6.
Two commensurable closed Fuchsian surfaces are tangent to the same discrete plane. The genus of a discrete plane $P$ is the smallest genus of the orientable Fuchsian surfaces tangent to $P$.
This is the genus $g_{c}(S)$ as in Theorem 1.8. Since there are only finitely many Fuchsian surfaces of genus up to $g$, and no distinct commensurability classes of Fuchsian surfaces are tangent, the set of planes up to genus $g$ is a disconnected closed set in $\mathcal{G}_{2}M$ (when this set has more than one planes). A subset $B$ of $\mathcal{G}_{2}M$ is filling if $\textnormal{Pr}(A)\subset M$ is filling. For $x\in M$ and $\Pi_{x}\in\mathcal{G}_{2}M$ tangent to a plane $P$, we call $\Pi_{x}$ intersects a geodesic $\gamma$ transversally in $M$ if $\gamma$ and $P$ intersect at some point transversally. Thus for a filling Fuchsian surface $S$, any $x\in S$, and any geodesic $\gamma\subset S$, $\Pi_{x}$ intersects $\gamma$ transversally by 3.18.
3.2 Large Fuchsian surfaces are strongly filling
One way to see that as the genus of discrete planes grows, they must cut $M$ into more and more components is as follows. Using 2.6, $M$ admits a fundamental domain $D\subset\mathbb{H}^{3}$, which is a compact convex polyhedron of diameter $d$. The intersection of $D$ with any geodesic plane in $\mathbb{H}^{3}$ has an area upper bound $2\pi(\cosh\frac{d}{2}-1)$ using the polar coordinates (1.3). Thus as the area of $S$ grows, the lift of $S$ in $D$ consists of more and more geodesic polygons in $D$. [MS95] implies that these polygons are uniformly distributed and thus must cut $M$ into components of decreasing diameter.
@@ 3.2 Large Fuchsian surfaces are strongly filling
$h_{2}(R)=\frac{\pi(1-\cos\theta_{0})}{4}(R\cosh R-R).$
@@
Let $s=s(z)$ be a unit-speed parameterization where $z\in I^{\prime}=[-R,R]$ and $s(I^{\prime})$ is embedded. Let $I=[\frac{-R}{2},\frac{R}{2}]$ parameterize the central half of $s$. Extend $z$ to the Fermi coordinates $(r,\theta,z)$ centered at $s$, where $r,\theta$ form the $2$-dimensional polar coordinate for the hyperbolic disk orthogonal to $s$ at $s(z)$. The hyperbolic metric is given by $\textnormal{d}r^{2}+\sinh^{2}r\,\textnormal{d}\theta^{2}+\cosh^{2}r\,\textnormal{d}z^{2}$. Let $\phi_{1},\phi_{2}$ be the spherical coordinates for the unit tangent space at $(r,\theta,z)$, where $\phi_{1}\in[0,\pi]$ measures the angle relative to $z$-axis and $\phi_{2}\in[0,2\pi]$ corresponds to $\theta$. Hence $(r,\theta,z,\phi_{1},\phi_{2})$ form local coordinates for $(x,v)$ in the unit tangent bundle $T^{1}M$ which double covers $\mathcal{G}_{2}M$.
We now describe an open subset $O(s)\subset\mathcal{G}_{2}M$ whose points are tangent to planes transversally intersecting $s(I)$ with angle $h_{1}(R)$. Using the ball model, we place the segment $s$ at a vertical $\mathbb{H}^{2}$ so that its perpendicular bisector passes through the origin. Then $s(I)=\{(0,0,z)|z\in[-R/2,R/2]\}$. The angle $\theta=0$ (resp. $\theta=\pi$) corresponds to the right (resp. left) half-space of the complement of the geodesic line containing $s(I)$ in $\mathbb{H}^{2}$. The coordinates $r\in[0,R/2]$ and $\theta\in[0,2\pi]$ parameterize a solid annulus $A$ centered at $s$. Let $\theta_{0}$ be the angle formed by the segment $\{(r,\pi,R/4)|r\in[0,R/2]\}$ and the segment $(R/2,\pi,R/4),(0,0,R/2)]$. By applying [Bus10, (i), (iii), (vi)] to the right triangle with side-length $a=R/4,b=R/2,\alpha=\theta_{0}$, and $\beta$, we see
@@
Recall $\mathbf{g}_{c}(S)$ denotes the minimal genus of an orientable surface commensurable with $S$. We first prove that the maximal diameter of components of $M-S_{i}$ tends to 0 as $g_{i}\coloneqq\mathbf{g}_{c}(S_{i})\rightarrow\infty$. Suppose, for contradiction, that there exists a subsequence $S_{i}$ with a component $C_{i}\subset M-S_{i}$ such that $\textnormal{diam}(C_{i})\geq 2R>0$ (we may assume $R<\textnormal{inj}(M)$).
Since $S_{i}$ is totally geodesic, each component of $\mathbb{H}^{3}-\pi^{-1}(S_{i})$ is the intersection of finitely many geodesic half-spaces and thus must be convex. In particular, $C_{i}$ contains a geodesic segment $\gamma_{i}$ of length $2R$. Since $M$ is compact, up to taking a subsequence, $\gamma_{i}$ converges to a geodesic segment $\gamma$ of length $2R$. Using 3.7, the subset of Grassmann bundles tangent to planes transversally intersecting the central half of $\gamma$ with an angle $\geq h_{1}(R)$ contains an open set $O$ of measure $\mu_{\mathcal{L}}(O)\geq h_{2}(R)>0$.
@@
Since the maximal diameter of components of $M-S_{i}$ tends to 0, their maximal volume also tends to 0. Thus the number of components necessarily tends to infinity. ∎
If a closed hyperbolic $n$-manifold $M^{n}$ contains a sequence of totally geodesic hypersurfaces $S_{i}$, since $S_{i}$ equidistribute by [MS95], the proof of Theorem 1.8 similarly shows that $S_{i}$ cuts $M^{n}$ into components whose diameters tend to 0. Each component $O$ of $M-S_{i}$ lifts to a convex subset of $\mathbb{H}^{n}$. If $\textnormal{diam}(O)<2\textnormal{inj}(M)$, then $O$ is contractible in $M$. This finishes the proof of the statement from Theorem 1.1: there exists a constant $A_{0}(M)$ such that every totally geodesic hypersurface whose Hausdorff measure $\geq A_{0}(M)$ is filling.
Given a closed hyperbolic $3$-manifold with exact generators $\{g_{1},\cdots,g_{m}\}$ of $\pi_{1}(M)\rightarrow\Gamma$ (in the sense of 2.6) for a fundamental domain $D$ with sides $F_{i},F_{i}^{\prime}$ (for $i=1,\cdots,m$), one can prove there exist $2m$ open subsets $O_{i},O_{i}^{\prime}$ of $\mathcal{G}_{2}M$ so that a Fuchsian $S$ whose $\mathcal{G}_{2}S$ passes through all $O_{i}$ and $O_{i}^{\prime}$ are filling, using 3.7. Each $O_{i}$ (resp. $O_{i}^{\prime}$) corresponds in $M$ to nearly orthogonal to $g_{i}$ and close to $F_{i}$ (resp. $F_{i}^{\prime}$) in the Grassmann bundle metric $d_{\mathcal{G}_{2}M}$.
We obtain the following corollary on the set $\mathcal{P}(M)$ of planes. Denote $\mathcal{P}_{\textnormal{fill}}(M)=\{P\in\mathcal{P}(M):M-P\textnormal{ consists of contractible components}\}.$
@@
The subset $\mathcal{P}_{\textnormal{fill}}(M)$ of filling planes is the dense open set $\mathcal{P}(M)-\sqcup_{i=1}^{m}P_{i}$ where $P_{i}$ corresponds to some non-filling surface.
@@
By Ratner’s [Rat91] and Shah’s [Sha91a] theorem, a plane corresponds to either a dense subset of $\mathcal{G}_{2}M$ or a closed Fuchsian surface. The projection of a dense subset of $\mathcal{G}_{2}M$ must be filling in $M$, and there are at most finitely many commensurability classes of non-filling Fuchsian surfaces by Theorem 1.1. ∎
@@
The above proof and hence Theorem 1.8 are valid for higher-dimensional hyperbolic manifolds and their totally geodesic hypersurfaces.
@@
Theorem 1.8 trivially holds for the second and third categories. [MR03, Theorem 9.5.4] shows how to construct infinitely many arithmetic hyperbolic $3$-manifolds containing Fuchsian surfaces. We first proved Theorem 1.8 and that large Fuchsian surfaces are filling. Vladimir Marković further suggested that most nearly geodesic surfaces Jeremy Kahn and he constructed in [KM12] should be filling, which led to part of the formulation of Theorem 1.1.

### v3

@@ 2.2 A natural metric on the Grassmann bundles
using parallel transport as follows. Given $\Pi\in(\mathcal{G}_{2})_{x}M$ and $\alpha:(-\eta,\eta)\rightarrow M$ such that $\alpha^{\prime}(0)\perp\Pi$ in $T_{x}M$, let $Z(t)$ be the parallel transport of $\Pi$ along $\alpha$. Let $\sigma:(-\eta,\eta)\rightarrow\mathcal{G}_{2}M$ be the curve
@@
There is an identification between $PSL_{2}(\mathbb{C})$ and the frame bundle $\mathcal{F}\mathbb{H}^{3}$ over $\mathbb{H}^{3}$. Moreover, $PSL_{2}(\mathbb{C})$ acts on $\mathcal{F}\mathbb{H}^{3}$ by isometries freely and transitively on the left. There is a right action on $\mathcal{F}\mathbb{H}^{3}$ by various flows, e.g., geodesic flow and unipotent flows. The bundle $\mathcal{G}_{2}\mathbb{H}^{3}$ is a natural quotient of $\mathcal{F}\mathbb{H}^{3}$: $\mathcal{G}_{2}\mathbb{H}^{3}=\mathcal{F}\mathbb{H}^{3}/SO(2)$. Each element in $\mathcal{G}_{2}\mathbb{H}^{3}$ is tangent to a unique geodesic plane $P$, an isometric embedding of $\mathbb{H}^{2}$ in $\mathbb{H}^{3}$. Since each $\mathbb{H}^{2}$ is a quotient of $PSL_{2}(\mathbb{R})$, by a $PSL_{2}(\mathbb{R})$-orbit in $\mathcal{G}_{2}\mathbb{H}^{3}$, we mean the projection of an orbit of $PSL_{2}(\mathbb{R})$ acting on $\mathcal{F}\mathbb{H}^{3}$ to $\mathcal{G}_{2}\mathbb{H}^{3}$. A lattice $\Gamma\leq PSL_{2}(\mathbb{C})$ acts on $\mathcal{F}\mathbb{H}^{3}$ or $\mathcal{G}_{2}\mathbb{H}^{3}$ on the left, and the orbits descend to $\mathcal{G}_{2}M\coloneqq\Gamma\backslash\mathcal{F}\mathbb{H}^{3}/SO(2)$.
2.3 Fundamental domains and fillingness
@@
By the first paragraph of [Rat19, 6.7], there exists a convex Dirichlet fundamental domain $D$. Furthermore, [Rat19, Theorem 6.7.1] proves that $\bar{D}$ is a convex polyhedron. Finally, [Rat19, Theorem 6.7.4(2), Theorem 6.7.5] prove that the fundamental domain $D$ is exact: for any side $F$ of $D$, there exists a unique $g\in\Gamma$ such that
@@
A closed hyperbolic $n$-manifold admits a fundamental domain $D\subset\mathbb{H}^{n}$ which is a convex polyhedron, whose sides are partitioned into pairs $\{\{F_{1},F_{1}^{\prime}\},\cdots,\{F_{k},F_{k}^{\prime}\}\}$, such that for each pair of sides $F_{i},F_{i}^{\prime}$, there exists a unique $g_{i}\in\Gamma$ which maps $F_{i}$ isometrically onto $F_{i}^{\prime}$. Moreover, the group $\Gamma$ is generated by $\{g_{1},\cdots,g_{k}\}$.
A surface representative $S\hookrightarrow M$ is filling if MSM-S consists of components that are contractible in $M$. A surface $S\hookrightarrow M$ is genuinely filling if MSM-S consists of topological balls. We always assume surfaces in hyperbolic $3$-manifolds are quasi-Fuchsian. Let $\pi:\mathbb{H}^{3}\rightarrow M$ be the covering projection. Then $\pi(\partial D)$ is a genuinely filling closed surface which is homotopically trivial. This motivates us to define a filling class of surfaces as a property that should hold for every representative. We thank Chao Wang for clarifications regarding filling surfaces versus filling homotopy classes. A filling class $[S]$ intersects every homotopically nontrivial loop of $M$ which means $\pi_{1}(M-S)\rightarrow\pi_{1}(M)$ is trivial. We record the following useful lemma from Rubinstein-Sageev.
@@
Let $M$ be a closed hyperbolic $3$-manifold and $[S]$ a filling class of surfaces. Then a least area representative $S$ is genuinely filling.
Thus, similar to the boundary of a fundamental domain, any surface in a filling class has an area lower bound by the linear isoperimetric inequality (see Theorem 1.11).
@@ 3 Strongly filling surfaces in hyperbolic $3$-manifolds
3.1 Some lemmas on intersections and fillingness
In this subsection, with a slight abuse of language, we also use $[S]$ to denote a homotopy class of surfaces (instead of modulo both homotopy and commensurability). We first define the intersection number between $\gamma]\in\pi_{1}(M)$ and a homotopy class $[S]$ of immersed surfaces. By applying homotopies, we may assume an immersed closed curve $\gamma:S^{1}\rightarrow M$ and an immersed surface $f:S\rightarrow M$ intersect at finitely many points $x_{1},\cdots,x_{n}$, such that for each $i\in\{1,\cdots,n\}$, there exist $j\in\{1,\cdots,I_{i}(\gamma)\}$, $k\in\{1,\cdots,I_{i}(S)\},t_{ij}\in S^{1}$, and $y_{ik}\in S$ with
@@ 3.1 Some lemmas on intersections and fillingness
$|\gamma\cap S|=\sum_{i=1}^{n}I_{i}(\gamma)\cdot I_{i}(S).$
@@
A Fuchsian surface minimizes the intersection number with a closed geodesic in its homotopy class.
@@
Lift $M$ to a fundamental domain $D\subset\mathbb{H}^{3}$. The preimages of $\gamma$ and $S$ in $D$ are finitely many geodesic arcs and geodesic polygons. If an arc which is a subset of a geodesic line $\tilde{\gamma}$ has a transverse intersection with a plane $P$, then the circle $\partial P$ separates the two endpoints of $\tilde{\gamma}$. By the Morse lemma, any homotopy of $\gamma$ induces a homotopy of $\tilde{\gamma}$, fixing its limit points. Similarly, any homotopy of $S$ induces a homotopy of $P$, fixing the circle at infinity. Thus no homotopy decreases the intersection number between $\beta$ and $P$. Applying this argument to all transverse intersections between the preimages of $\gamma$ and the preimages of $S$ through $D$, we conclude that the minimal number is realized by taking the geodesic representatives of both $\gamma$ and $S$. ∎
We call the intersection between a Fuchsian surface and a closed geodesic in a hyperbolic manifold essential. The strong interplay between topology and geometry is a recurring theme in hyperbolic geometry. If $M=S^{2}\times S^{1}$ and $\gamma\subset S^{2}$ a geodesic great circle, then $S\coloneqq\gamma\times S^{1}\subset M$ is a compressible totally geodesic torus. Thus for another geodesic $\eta$, $\eta]\cdot[S]<|\eta\cap S|$ and 3.2 fails in general.
@@
If a Fuchsian surface $S$ intersects a geodesic $\gamma$ transversally, then for any $S^{\prime}$ homotopic to $S$, $\pi_{1}(M-S^{\prime})$ does not contain $\gamma]$. Moreover, if a Fuchsian surface $S$ is genuinely filling, its homotopy class and commensurability class are strongly filling.
@@
3.2 and 3.3 do not generalize to least area non-Fuchsian surfaces in hyperbolic manifolds: 3.21 proves every such surface has redundant intersections with some closed geodesics.
@@
Let $S\hookrightarrow M$ be a filling surface and $M^{\prime}$ a cover of $M$. Then the preimage of $S$ in $M^{\prime}$ is a (possibly disconnected or noncompact) filling surface $S^{\prime}\subset M^{\prime}$.
@@
$\mathcal{P}(M)\coloneqq\mathcal{G}_{2}M/\sim$ (3.1)
where $(x_{1},\Pi_{1})\sim(x_{2},\Pi_{2})$ if they are belong to the same $PSL_{2}(\mathbb{R})$-orbit in $\mathcal{G}_{2}M$. This is a three-dimensional non Hausdorff space. By a slight abuse of language, we identify a plane with the subset it is tangent to in $\mathcal{G}_{2}M$. If $\Pi_{x}$ is a $2$-plane tangent at $x$, then $P(\Pi_{x})$ is the plane in $\mathcal{G}_{2}M$ passing through $\Pi_{x}$. By identifying the frame bundle of $\mathbb{H}^{3}$ as $PSL_{2}(\mathbb{C})$, we identify $\mathcal{P}(\mathbb{H}^{3})\cong PSL_{2}(\mathbb{C})/PSL_{2}(\mathbb{R})$ and $\mathcal{P}(M)=\Gamma\backslash PSL_{2}(\mathbb{C})/PSL_{2}(\mathbb{R})$. The projection map $\mathcal{G}_{2}M\rightarrow\mathcal{P}(M)$ is continuous, and the fiber of the projection map is isomorphic to $\mathbb{H}^{2}$.
For convenience later, if a plane $P$ corresponds to a closed Fuchsian surface $S$, we call $P$ discrete. The genus of a discrete plane is then defined as the genus $\mathfrak{g}[S]$. Since there are only finitely many Fuchsian surfaces of genus up to $g$, and no distinct commensurability classes of Fuchsian surfaces are tangent, the set of planes up to genus $g$ is a disconnected closed set in $\mathcal{G}_{2}M$ (when this set has more than one planes). A subset $B$ of $\mathcal{G}_{2}M$ is filling if $\textnormal{Pr}(B)\subset M$ is filling. For $x\in M$ and $\Pi_{x}\in\mathcal{G}_{2}M$ tangent to a plane $P$, we say $\Pi_{x}$ intersects a geodesic $\gamma$ transversally in $M$ if $\gamma$ and $P$ intersect at some point transversally. Thus for a filling Fuchsian surface $S$, any $x\in S$, and any geodesic $\gamma\subset S$, $\Pi_{x}$ intersects $\gamma$ transversally by 3.6.
3.2 A filling totally geodesic surface is strongly filling
Recall that a homotopy class $[S]$ of surfaces in a hyperbolic manifold $M$ separates $p,q\in\partial\widetilde{M}$ if any $S^{\prime}\in[S]$ has a cover $\widetilde{S}$ such that $\partial\widetilde{S}$ separate $p,q\in\partial\widetilde{M}$. Contractible surfaces in $M$ can be homotoped to be filling. However, [RS99, Lemma 2.1] proved that if a Fuchsian surface $S$ is filling, its class $[S]$ is strongly filling. We present a slightly simpler proof.
Lemma 3.6.
Proof.
@@ 3.2 A filling totally geodesic surface is strongly filling
and $\partial B_{i}=\cup_{j=1}^{m}F_{j}$ where $F_{j}$ are geodesic polyhedrons. Since $B_{i}$ is contractible in $M$, $B_{i}$ lifts isometrically to $\mathbb{H}^{n}$, which implies that $B_{i}$ is convex, and hence contractible as a topological space itself. Thus $S$ is genuinely filling.
[…]
Thus $\gamma\cap\partial B_{i}\neq\emptyset$ for some $i$. If $\gamma$ transversally intersects the interior of some polyhedron $F_{j}$ transversally, by lifting $\gamma$ and $F_{j}$ to $\mathbb{H}^{n}$, we find a plane tangent to $\widetilde{F_{j}}$ which separates $\partial\tilde{\gamma}$.
Corollary 3.7.
Let $M$ be a closed hyperbolic $n$-manifold. A totally geodesic hypersurface $S\hookrightarrow M$ is filling if and only if every component of $\widetilde{M}-\pi^{-1}(S)$ is bounded.
Remark 3.8.
The argument above relies on the analysis of the intersection of a geodesic $\gamma$ and a geodesic face $\subset$ a plane $P$ and the dichotomy that either $\tilde{\gamma}$ is a subset of $\widetilde{P}$ or $\tilde{\gamma}$ transversally intersects $\widetilde{P}$. This gives a geometric criterion for a homotopy class of totally geodesic surfaces to be filling in certain aspherical manifolds, see Section 4.1 for Euclidean and $\mathbb{H}^{2}\times\mathbb{R}$ 33-manifolds.
3.3 All but finitely many totally geodesic hypersurfaces are strongly filling
One way to see that as the genus of discrete planes grows, they must cut $M$ into more and more components is as follows. Using 2.6, $M$ admits a fundamental domain $D\subset\mathbb{H}^{3}$, which is a compact convex polyhedron of diameter $d$. The intersection of $D$ with any plane in $\mathbb{H}^{3}$ has an area upper bound $2\pi(\cosh\frac{d}{2}-1)$ using the polar coordinates (1.3). As the area of $S$ grows, the preimage of $S$ in $D$ consists of more and more geodesic polygons in $D$. [MS95] implies that these polygons are uniformly distributed and thus must cut $M$ into components of decreasing diameter.
@@ 3.3 All but finitely many totally geodesic hypersurfaces are strongly filling
$h_{2}(R)=\frac{\pi(1-\cos\theta_{0})}{4}(R\cosh R-R),$
where $\theta_{0}=\cot^{-1}\left(2\cosh^{2}\frac{R}{4}\right)$. The following lemma works similarly in every dimension. For the clarity of presentation, we fix dimension $n=3$.
@@
Let $s=s(z)$ be a unit-speed parameterization where $z\in I^{\prime}=[-R,R]$ and $s(I^{\prime})$ is embedded. Let $I=[\frac{-R}{2},\frac{R}{2}]$ parameterize the central half of $s$. Extend $z$ to the cylindrical coordinates $(r,\theta,z)$ centered at $s$, where $r,\theta$ form the $2$-dimensional polar coordinate for the hyperbolic disk orthogonal to $s$ at $s(z)$. The hyperbolic metric is given by $\textnormal{d}r^{2}+\sinh^{2}r\,\textnormal{d}\theta^{2}+\cosh^{2}r\,\textnormal{d}z^{2}$ (see [BD17, p. 548]), where $r=0$ corresponds to the geodesic $s$, and $r=\text{constant}$ gives concentric Euclidean annulus. Let $\phi_{1},\phi_{2}$ be the spherical coordinates for the unit tangent space at $(r,\theta,z)$, where $\phi_{1}\in[0,\pi]$ measures the angle relative to $z$-axis and $\phi_{2}\in[0,2\pi]$ corresponds to $\theta$. Hence $(r,\theta,z,\phi_{1},\phi_{2})$ form local coordinates for $(x,v)$ in the unit tangent bundle $T^{1}M$ which double covers $\mathcal{G}_{2}M$.
We now describe an open subset $O(s)\subset\mathcal{G}_{2}M$ whose points are tangent to planes transversally intersecting $s(I)$ with angle $h_{1}(R)$. Using the ball model, we place the segment $s$ at a vertical $\mathbb{H}^{2}$ so that its perpendicular bisector passes through the origin. Then $s(I)=\{(0,0,z)|z\in[-R/2,R/2]\}$. The angle $\theta=0$ (resp. $\theta=\pi$) corresponds to the right (resp. left) half-space of the complement of the geodesic line containing $s(I)$ in $\mathbb{H}^{2}$. The coordinates $r\in[0,R/2]$ and $\theta\in[0,2\pi]$ parameterize a solid cylinder $A$ centered at $s$. Let $\theta_{0}$ be the angle formed by the segment $\{(r,\pi,R/4)|r\in[0,R/2]\}$ and the segment $(R/2,\pi,R/4),(0,0,R/2)]$. Combining [Bus10, (i), (iii), (vi)], we deduce that for a right triangle with legs $a$ and $b$, opposite to angles $\alpha$ and $\beta$, respectively, we have
$\frac{\cos\alpha}{\sin\alpha}=\frac{\sinh b\cosh a}{\sinh a}.$
Set $a=R/4,b=R/2,\alpha=\theta_{0}$, we see
@@
We give a proof for $n=3$ and for higher dimensions the proof works similarly. We first prove that the maximal diameter of components of $M-S_{i}$ tends to 0 as $g_{i}\coloneqq\mathfrak{g}[S_{i}]\rightarrow\infty$. Suppose, for contradiction, that there exists a subsequence $S_{i}$ with a component $C_{i}\subset M-S_{i}$ such that $\textnormal{diam}(C_{i})\geq 2R>0$ (we may assume $R<\textnormal{inj}(M)$).
Since $S_{i}$ is totally geodesic, each component of $\mathbb{H}^{3}-\pi^{-1}(S_{i})$ is the intersection of finitely many geodesic half-spaces and thus must be convex. In particular, $C_{i}$ contains a geodesic segment $\gamma_{i}$ of length $2R$. Since $M$ is compact, up to taking a subsequence, $\gamma_{i}$ converges to a geodesic segment $\gamma$ of length $2R$. Using 3.9, the subset of Grassmann bundles tangent to planes transversally intersecting the central half of $\gamma$ with an angle $\geq h_{1}(R)$ contains an open set $O$ of measure $\mu_{\mathcal{L}}(O)\geq h_{2}(R)>0$.
@@
Since the maximal diameter of components of $M-S_{i}$ tends to 0, their maximal volume also tends to 0. Thus the number of components necessarily tends to infinity. If a component of $M-S_{i}$ has diameter $\textnormal{inj}(M)$, it is contractible. ∎
Given a closed hyperbolic $n$-manifold with the side-pairings $\{g_{1},\cdots,g_{k}\}$ of $\pi_{1}(M)\rightarrow\Gamma$ (in the sense of 2.6) for a fundamental domain $D$ with sides $F_{i},F_{i}^{\prime}$ (for $i=1,\cdots,k$), one can prove there exist $2k$ open subsets $O_{i},O_{i}^{\prime}$ of $\mathcal{G}_{n-1}M$ so that a totally geodesic hypersurface $S$ whose $\mathcal{G}_{n-1}M$ passes through all $O_{i}$ and $O_{i}^{\prime}$ are filling, using 3.9. Each $O_{i}$ (resp. $O_{i}^{\prime}$) is a subset of $\mathcal{G}_{n-1}M$ tangent to planes nearly orthogonal to the axis of $g_{i}$ and based at points close to $F_{i}$ (resp. $F_{i}^{\prime}$).
Let $\mathcal{P}(M)$ be the set of planes.
@@
Let $M$ be a closed hyperbolic $n$-manifold. All but finitely many planes in $\mathcal{P}(M)$ are strongly filling.
@@
By Ratner’s [Rat91] and Shah’s [Sha91a] theorem, a plane corresponds to either a dense subset of $\mathcal{G}_{n-1}M$ or a closed totally geodesic hypersurface. The projection of a dense plane of $\mathcal{G}_{n-1}M$ must be strongly filling in $M$, and there are at most finitely many non-filling discrete planes in $M$ by Theorem 1.6. ∎
@@
Theorem 1.6 trivially holds for the second and third categories. [MR03, Theorem 9.5.4] shows how to construct infinitely many arithmetic hyperbolic $3$-manifolds containing Fuchsian surfaces.